OpenStudy (anonymous):

A large diamond with a mass of 712.04 grams was recently discovered in a mine. If the density of the diamond is 5.02 g/cm^3 , what is the volume? Round your answer to the nearest hundredth. (5 points) Select one: a. 141.84 cm3 b. 507.02 cm3 c. 717.06 cm3 d. 3574.44 cm3

OpenStudy (anonymous):

@ganeshie8

OpenStudy (gudden):

D=M/V where d = density m = mass and V = volume..

OpenStudy (gudden):

@prowrestler ... now... use the formula.. that i have given by writing the values that the question has given for the 2 known...

OpenStudy (anonymous):

mass is 712.04 grams

OpenStudy (anonymous):

density of the diamond is 5.02 g/cm^3

OpenStudy (gudden):

so @prowrestler .. use those values in the equation that i have given u...

OpenStudy (anonymous):

what do i do whith those numbers do i plug them in

OpenStudy (anonymous):

You may notice that density is measured in g/cm3. Thus when density is 5.02 g/cm^3, you could also say that "it has 5.02g per 1 cm3", or "each cm3 contains 5.02 grams". Then you need to ask yourself, if 1cm3 can contain a mass of 5.02g, how many cm3 would be need to put in 712.04 grams of mass. Think of the cubic centimetre like a basket that you put weight in. How many would be needed?

OpenStudy (gudden):

yeah @prowrestler .. im solving it here u can take a look D = M/V i.e. V = M/D so just divide the mass given i.e.712.04 grams with 5.02 and u would get the answer

OpenStudy (anonymous):

712.04/5.02

OpenStudy (gudden):

yeah @prowrestler ...:D

OpenStudy (anonymous):

i got 141.84063..... is that correct

OpenStudy (anonymous):

@gudden @ganeshie8

OpenStudy (gudden):

yeah u r correct

OpenStudy (anonymous):

thanks for all the help

OpenStudy (anonymous):

can you come help me with some more

OpenStudy (gudden):

sure...